A round swimming pool has a diameter of 20 feet and is 4 feet tall. How much water will the pool hold?
Found 2 solutions by c-dubb swimmer, checkley77:
Answer by c-dubb swimmer(3)

formula for volume of a cylinder: Pi times Radius^2 (squared) time height
(100)(4)= 400Pi (for this problem i used Pi = 3.14)
400 times 3.14 = 1256
The pool will hold about 1256 (insert unit [gallons, liters, etc]).

Answer by checkley77(12844)

pi*r^2*h=3.14*10^2*4=1,256 ft.^3 is the volume of the pool.